22west:

1. Refer to Expression 1 and Expression 2. Expression 1 Expression 2 (a) What is the value of Expression 1? (b) Insert parentheses in Expression 2 so that it has a value of 19. Then show why your expression has a value of 19.

22west:

Equation 1: \[6\times7-3^{2}\times9+4^{3}\]

22west:

Equation 2: \[5+4\times2+6-2\times2-1\]

22west:

@dude

ZoeyBiocth16:

6*7-3^2*9+4^3= 42-3^2*9+4^3 42-9*9+64 25>>>i hope its right 5+4×2+6−2×2−1=14

22west:

Did you read the question?

22west:

(b) Insert parentheses in Expression 2 so that it has a value of 19. Then show why your expression has a value of 19.

ZoeyBiocth16:

*face palm*

22west:

Lol

22west:

I put the parenthesis around 5+4 and got 19 but my teacher marked it incorrect because I didn't show why... >:(

ZoeyBiocth16:

(5+(4*2)+6)-(2*2)-1

ZoeyBiocth16:

) at the end of the 2

dude:

Zoey use \(\LaTeX\) ;p \(6\times 7-3^2\times 9+4^3=\\ 6\times 7-9\times 9+64\\ 42-81+64\\ 25\)

ZoeyBiocth16:

yea dude i dont know how but imma go see

22west:

Okay so Equation 1 equals 25. What about the second part of the question?

ZoeyBiocth16:

i gtg eat

22west:

(b) Insert parentheses in Expression 2 so that it has a value of 19. Then show why your expression has a value of 19.

dude:

For the parenthesis between 5 and 4 you were right (Originally you had to solve using PEMDAS) To explain that: \((5+4)×2+6−2×2−1\) Once you have parenthesis you have to solve 5+4 first (PEMDAS) \(9\times 2+6-2\times 2-1\) Then Multiply (PEMDAS) \(18+6-4-1\) Simplify \(=19\)

22west:

Thank you sir